OpenStudy (silverlark2):

Coordinates of two points are A (-2,6) and B (9.3). Find the coordinates of a point C on the x axis such that AC = BC.

OpenStudy (anonymous):

call that point (x,0) second coordinate is 0 becaues you are on the x axis

OpenStudy (anonymous):

then you know that \[(x+2)^2+6^2=(x-9)^2+3^2\] by the distance formula. now you have a quadratic equation to solve

OpenStudy (silverlark2):

ty :)

OpenStudy (anonymous):

actually i lied, you do not have a quadratic equation to solve, because when you expand the x^2 term will drop out

OpenStudy (anonymous):

\[x^2+4 x+40 = x^2-18 x+90\] then \[4x+40=-18x+90\] etc

OpenStudy (silverlark2):

Oh haha ok then. Thanks again though